# Cross-Camera Distracted Driver Classification through Feature Disentanglement and Contrastive Learning

### v1

@@ Abstract
The classification of distracted drivers is pivotal for ensuring safe driving. Previous studies demonstrated the effectiveness of neural networks in automatically predicting driver distraction, fatigue, and potential hazards. However, recent research has uncovered a significant loss of accuracy in these models when applied to samples acquired under conditions that differ from the training data. In this paper, we introduce a robust model designed to withstand changes in camera position within the vehicle. Our Driver Behavior Monitoring Network (DBMNet) relies on a lightweight backbone and integrates a disentanglement module to discard camera view information from features, coupled with contrastive learning to enhance the encoding of various driver actions. Experiments conducted on the daytime and nighttime subsets of the 100-Driver dataset validate the effectiveness of our approach with an increment on average of 9% in Top-1 accuracy in comparison with the state of the art. In addition, cross-dataset and cross-camera experiments conducted on three benchmark datasets, namely AUCDD-V1, EZZ2021 and SFD, demonstrate the superior generalization capability of the proposed method.
@@ I Introduction
As of 2019, road traffic accidents are the leading killer of children and youth aged 5 to 29 years and are the 12th leading cause of death when all ages are considered [1]. Traffic accidents pose a growing problem and are projected to rank as the seventh leading cause of global mortality by 2030.111https://www.who.int/data/gho/data/themes/topics/topic-details/GHO/road-traffic-mortality (accessed: Jul 31, 2024). According to statistics from the European Road Safety Observatory (ERSO), 5 to 25% of traffic accidents in Europe stem from driver distraction. [^2] https://road-safety.transport.ec.europa.eu/system/files/2021-07/ersosynthesis2018-driverdistraction-summary.pdf (accessed: Jul 31, 2024). The ERSO defines distracted driving as “the diversion of attention away from activities critical for safe driving toward a competing activity, which may result in insufficient or no attention to activities critical for safe driving”. Commonly observed distractions include smartphone usage, eating, or conversing with passengers, all of which heighten the risk of accidents and should be rigorously avoided. Detecting these distracting behaviors through automatic in-vehicle computational systems enables early intervention to avoid distraction-related accidents. In this context, automatic Driver Monitoring Systems (DMS) have emerged as a novel Advanced Driver Assistance System (ADAS) technology to mitigate car accidents [2]. DMSs exploit machine and deep learning techniques and are trained on datasets acquired in both real and simulated scenarios, often with varying protocols and experimental setups [3, 4, 5, 6]. Significant differences among these setups include the interior configuration of the vehicle, the position and posture of the driver and the camera placement. Consequently, models trained on specific datasets exhibit poor generalization to different data [7]. To deploy these models on non-production embedded systems in various vehicles while maintaining performance, the models must be robust to prior variations. Furthermore, they must generalize to new conditions without increasing computational complexity, considering the limited capacity of the target devices.
@@
This paper addresses the challenge of cross-camera distracted driver classification by introducing the Driver Behavior Monitoring Network (DBMNet). This model is based on the use of a lightweight Convolutional Neural Network (CNN) for encoding RGB images to facilitate its deployment on commercial off-the-shelf devices. To make the model robust across different camera views, we propose a feature disentanglement module. This module separates action-relevant features from view-relevant features, allowing the proposed model to learn high quality features that are invariant to camera and view variations. This is achieved by a weighted combination of the view queries and the feature vector determined by the backbone. The weights are calculated on the basis of the similarity between the test sample and the training views. Moreover, we employ a supervised contrastive learning approach with triplet loss to enhance the learning of disentangled features. While cross-entropy loss theoretically allows the model to discern similarities within categories and differences between them, it does not explicitly guarantee this outcome. In contrast, triplet loss explicitly compels the model to identify similar features for samples within the same category and distinguish features for samples from different categories. In our model, one triplet loss brings representations of different actions captured by the same camera closer together while separating them from identical actions captured by the same camera. This improves the model’s ability to discriminate between different views and enhances the effectiveness of the disentanglement module. Another triplet loss brings representations of identical actions captured by different cameras closer together while separating different actions captured by the same camera. This aids in the categorization of driver actions.
@@
The proposed work is unique because no previous approach has explicitly addressed modeling invariance in experimental configurations. We adopt a leave-one-camera-out approach, which mirrors the realistic scenario where abundant data from various cameras is available for training, and experimentation occurs under diverse conditions. We demonstrate that our novel network and learning procedure alleviates the degradation in cross-camera distracted driver classification on the very recent 100-Driver dataset [4]. Experiments on the AUCDD-V1 [3], EZZ2021 [8], and SFD [9] datasets demonstrate the superior cross-dataset and cross-camera performance of our approach, highlighting its generalization capability. Our main contributions can be summarized as follows:
@@
  • Two triplet losses to aid in learning disentangled representations.
@@ II Related works
Researchers simplified training strategies for enhanced generalization. Masood et al. [28] used pretrained-weight VGG architectures for reduced training time. Li et al. [29] introduced Multi-Teacher Knowledge Distillation for lightweight models. Peng et al. [30] employed Few-Shot Learning for feature calibration. Contrastive and unsupervised methods, like Hu et al. [5] and Li et al. [31], focused on improved training strategies.
Compared to previous research, this paper tackles the issue of distracted drivers by enhancing model robustness without increasing computational demands. Notably, this is the first work to conduct comprehensive experiments using the leave-one-camera-out strategy, accurately reflecting real-world scenarios where the recognition model is trained under certain conditions and tested under different ones. This rigorous setup ensures our findings are both practical and applicable to real-world conditions.
@@ III-A Motivation
Image analysis-based DMSs encounter a significant challenge due to the variability in vehicle interior configurations, resulting in disparate camera viewpoints [30, 4]. Conventional supervised deep learning methods, often employing cross-entropy loss, struggle to develop robust representations capable of generalizing across different views. In these varied perspectives, where driver actions are observed, it becomes crucial to design supervised techniques that mitigate the drop in effectiveness when handling data acquired under conditions that differ from the training set.
Moreover, deploying these DMSs on devices with limited computational capacity necessitates the use of efficient and lightweight models, ensuring real-time predictions. Hence, robust solutions must be sought, incorporating deep learning models that respond to previous requirements.
@@ III-B Overview
Given the factors mentioned above, we present the Driver Behavior Monitoring Network (DBMNet), illustrated in Figure 1. This method encodes an RGB image by leveraging the lightweight CNN architecture detailed in [4] as its backbone. The proposed feature disentanglement module aims to prioritize action encoding over camera-view information. This is achieved by linearly combining view queries with features estimated by the backbone. The resulting disentangled features are then forwarded to a linear layer for action classification. The model outputs two probability distributions: one for the driver’s action categories and another for the viewpoint. The viewpoint distribution is used solely as an auxiliary task for the feature disentanglement module. The model is optimized through cross-entropy losses combined with triplet losses to assist in the learning of view-invariant action features.
Figure 1: The DBMNet pipeline. At inference time (a), the backbone encodes an RGB image into a feature vector $\mathbf{f}$. The feature disentanglement module then refines these features to discard view-related information while retaining action-related details using view queries, Q, together with the probability distribution over the views. The resulting features $\mathbf{\hat{f}}$ are mapped into a driver action by the action classifier. During training (b), alongside the cross-entropy loss for action and view prediction (respectively, $\mathcal{L}_{ace}$ and $\mathcal{L}_{vce}$), we employ the triplet loss ($\mathcal{L}_{ac}$ and $\mathcal{L}_{vc}$). This triplet loss helps to learn distinct action and view representations by processing three input images: an anchor image $a$, an image $sv$ depicting the same view but different action as $a$, and an image $sa$ with the same action but from a different view as $a$.
@@ III-C The proposed model
where $\mathbf{p}_{v}$ is the probability distribution over $V$ views that is estimated by the view classifier.
@@ III-D Training procedure
$\mathcal{L}=\mathcal{L}_{ace}+\mathcal{L}_{vce}+\mathcal{L}_{ac}+\mathcal{L}_{vc}.$ (5)
In the previous loss, there are two cross-entropy losses mainly dedicated to the optimization of linear layers for view ($\mathcal{L}_{vce}$) and action ($\mathcal{L}_{ace}$) categorization. Cross-entropy losses are only applied to the anchor image in this way:
$\mathcal{L}_{vce}=-\sum_{m=1}^{V}y^{(m)}_{v}\mathrm{log}(p_{v}^{(m)}),$ (6)
$\mathcal{L}_{ace}=-\sum_{m=1}^{A}y^{(m)}_{a}\mathrm{log}(p_{a}^{(m)}).$ (7)
@@
The previous two cross-entropy losses are complemented by two opposite triplet losses. Both losses use the same triplet construction previously defined. The difference lies in the role assumed by $sv$ ed $sa$ in each loss.
In the action triplet loss (Eq. 8) the objective is to learn features for action encoding that are invariant regardless of the viewpoint. Hence, $sa$ serves as the positive sample, whereas $sv$ serves as the negative sample. This choice has the purpose of bringing closer the representations of $sa$ and $a$ which have the same action. Conversely, the distance between the representations of $a$ and $sv$ that depict a different action from the same point of view is instead maximized:
$\mathcal{L}_{ac}=\sum_{n=1}^{N}[\delta+\mathcal{D}(\mathbf{\hat{f}}_{a},\mathbf{\hat{f}}_{sa})-\mathcal{D}(\mathbf{\hat{f}}_{a},\mathbf{\hat{f}}_{sv})],$ (8)
@@
For the view triplet loss, positive and negative samples are reversed:
$\mathcal{L}_{vc}=\sum_{n=1}^{N}[\delta+\mathcal{D}(\mathbf{f}_{a},\mathbf{f}_{sv})-\mathcal{D}(\mathbf{f}_{a},\mathbf{f}_{sa})].$ (9)
@@ 5 Texting (left) 11 Looking down 17 Adjusting radio
The driving sessions were recorded in both daytime and nighttime in five different vehicles, comprising two sedans, two SUVs, and one van. Within each vehicle, four cameras were strategically placed - on the left front, right front, and right side of the drivers. This arrangement served dual purposes. Firstly, having a dataset with multiple cameras aids in developing more robust models that can account for variations in camera perspectives, thus evaluating the models’ generalization across different cameras. Secondly, the multi-camera dataset offers an opportunity to enhance system performance by considering the content captured from multiple camera angles. Figure 2 reports samples from the four cameras in both day and night times.

### v2

@@ Abstract
The classification of distracted drivers is pivotal for ensuring safe driving. Previous studies demonstrated the effectiveness of neural networks in automatically predicting driver distraction, fatigue, and potential hazards. However, recent research has uncovered a significant loss of accuracy in these models when applied to samples acquired under conditions that differ from the training data. In this paper, we introduce a robust model designed to withstand changes in camera position within the vehicle. Our Driver Behavior Monitoring Network (DBMNet) relies on a lightweight backbone and integrates a disentanglement module to discard camera view information from features, coupled with contrastive learning to enhance the encoding of various driver actions. Experiments conducted using a leave-one-camera-out protocol on the daytime and nighttime subsets of the 100-Driver dataset validate the effectiveness of our approach. Cross-dataset and cross-camera experiments conducted on three benchmark datasets, namely AUCDD-V1, EZZ2021 and SFD, demonstrate the superior generalization capabilities of the proposed method. Overall DBMNet achieves an improvement of 7% in Top-1 accuracy compared to existing approaches. Moreover, a quantized version of the DBMNet and all considered methods has been deployed on a Coral Dev Board board. In this deployment scenario, DBMNet outperforms alternatives, achieving the lowest average error while maintaining a compact model size, low memory footprint, fast inference time, and minimal power consumption.
@@ I Introduction
As of 2019, road traffic accidents are the leading killer of children and youth aged 5 to 29 years and are the 12th leading cause of death when all ages are considered [1]. Traffic accidents pose a growing problem and are projected to rank as the seventh leading cause of global mortality by 2030 [2]. According to statistics from the European Road Safety Observatory (ERSO), 5 to 25% of traffic accidents in Europe stem from driver distraction [3]. The ERSO defines distracted driving as “the diversion of attention away from activities critical for safe driving toward a competing activity, which may result in insufficient or no attention to activities critical for safe driving”. Commonly observed distractions include smartphone usage, eating, or conversing with passengers, all of which heighten the risk of accidents and should be rigorously avoided. Detecting these distracting behaviors through automatic in-vehicle computational systems enables early intervention to avoid distraction-related accidents. In this context, automatic Driver Monitoring Systems (DMS) have emerged as a novel Advanced Driver Assistance System (ADAS) technology to mitigate car accidents [4]. Recent legislative developments in the European Union further underline the relevance of this research direction. Specifically, Regulation (EU) 2019/2144 and the more recent Commission Delegated Regulation (EU) 2023/2590 mandate the adoption of driver drowsiness and distraction monitoring systems in all new vehicles starting from July 2024 and July 2026, respectively [5, 6]. DMSs exploit machine learning and deep learning techniques and are trained on datasets acquired in both real and simulated scenarios, often with varying protocols and experimental setups [7, 8, 9, 10]. Significant differences among these setups include the interior configuration of the vehicle, the position and posture of the driver, and the camera placement. Consequently, models trained on specific datasets exhibit poor generalization to different ones [11]. To deploy these models on non-production embedded systems in various vehicles while maintaining performance, the models must be robust to prior variations. Furthermore, they must generalize to new conditions without increasing computational complexity, considering the limited capacity of the target devices.
@@
This paper addresses the challenge of cross-camera distracted driver classification by introducing the Driver Behavior Monitoring Network (DBMNet). This model is based on the use of a lightweight Convolutional Neural Network (CNN) for encoding RGB images to facilitate its deployment on commercial off-the-shelf devices. To make the model robust across different camera views, we propose a feature disentanglement module. This module separates action-relevant features from view-relevant features, allowing the proposed model to learn high quality features that are invariant to camera and view variations. To achieve this, the module is designed as a hyper-network to compute a weighted combination of the view queries and the feature vector determined by the backbone. The weights are calculated on the basis of the similarity between the test sample and the training views. Moreover, we employ a supervised contrastive learning approach with triplet loss to enhance the learning of disentangled features. While cross-entropy loss theoretically allows the model to discern similarities within categories and differences between them, it does not explicitly guarantee this outcome. In contrast, triplet loss explicitly compels the model to identify similar features for samples within the same category and distinguish features for samples from different categories. In our model, one triplet loss brings representations of different actions captured by the same camera closer together while separating them from identical actions captured by the same camera. This improves the model’s ability to discriminate between different views and enhances the effectiveness of the disentanglement module. A second triplet loss brings representations of identical actions captured by different cameras closer together while separating different actions captured by the same camera. This aids in the categorization of driver actions.
@@
The proposed work is unique because no previous approach has explicitly addressed modeling invariance in experimental configurations. We adopt a leave-one-camera-out approach, which mirrors the realistic scenario where abundant data from various cameras is available for training, and experimentation occurs under diverse conditions. We demonstrate that our novel network and learning procedure alleviates the degradation in cross-camera distracted driver classification on the very recent 100-Driver dataset [8]. Experiments on the AUCDD-V1 [7], EZZ2021 [12], and SFD [13] datasets demonstrate the superior cross-dataset and cross-camera performance of our approach, highlighting its generalization capability. To the best of our knowledge, this is the first work to explicitly address cross-camera generalization in distracted driver classification through a leave-one-camera-out protocol on the 100-Driver dataset—a setup that closely mirrors real-world deployment scenarios and remains largely unexplored in prior research. Our main contributions can be summarized as follows:
@@
  • Two triplet losses to explicitly aid in learning disentangled representations.
@@
  • An evaluation of quantized models on a Coral Dev Board, demonstrating suitability for on-edge deployment in embedded automotive systems.
@@ II Related works
Researchers simplified training strategies for enhanced generalization. Masood et al. [32] used pretrained-weight VGG architectures for reduced training time. Li et al. [33] introduced Multi-Teacher Knowledge Distillation for lightweight models. Peng et al. [34] employed Few-Shot Learning for feature calibration. Recently, several approaches have leveraged self-attention mechanisms in Vision Transformers (ViTs) to address driver monitoring tasks [35, 36]. Mohammed et al. [35] proposed a semi-supervised lightweight hybrid ViT for accurate and efficient detection of driver distraction behaviors. DGCCL [36] utilized CLIP’s image encoder for feature extraction to leverage its strong generalization capabilities. To address domain shift, it introduced center loss, and further improved cross-domain performance by integrating AAMP into the classification loss. Finally, unsupervised methods like Li et al. [37], focused on improved training strategies.
In order to learn more general feature representations that can later be used for classification or clustering, researchers have investigated contrastive learning exploiting contrastive and triplet loss. A first group of contrastive loss-based papers uses supervised contrastive loss to align representations of the same class (i.e., action) and separate those of different classes. For example, Kopuklu et al. [38] proposed a contrastive learning approach to learn a metric to differentiate normal driving from anomalous driving and introduced a new Driver Anomaly Detection (DAD) dataset with front and top drivers views in both infrared and depth modalities. Koay et al. [39] proposed a deep contrastive learning approach exploiting multi-view and multi-modal data of DAD dataset to identify normal and anomalous driving in an open set manner, fusing together the results of different views and modalities.
More often, instead, they rely on self-supervised contrastive learning to learn generalizable embeddings without explicit labels for every behavior and typically use temporal or multimodal consistency, or augmentations, to form positive and negative pairs. Yang et al. [40] proposed a weakly supervised contrastive learning framework, in which distracted behaviors are identified based on their distances from the normal driving representation set. Hu et al. [9] extended the previous approach by proposing a novel clustering supervised contrastive loss to optimize the distribution of the extracted representation vectors, so that normal representations are better clustered while abnormal ones are separated. Khan et al. [41] presented a new supervised contrastive loss function along with considerations for including projection head and refining labels to improve the detection of anomalous driving behaviors from videos.
Triplet loss-based papers instead construct triplets (anchor, positive, negative) to ensure that feature embeddings of similar driver actions (i.e., same class) are closer and those of different actions are farther apart. Okon et al. [42] used the triplet loss with an offline selection of hard triplets to finetune a neural network pre-trained with classification loss. Liu et al. [43] used the triplet loss in a multi-task learning strategy to force the networks to explore global information by multiple tasks. The negative sample in each triplet is generated by randomly shuffling the local regions of each given input.
Compared to previous research, this paper tackles the issue of distracted drivers by enhancing model robustness without increasing computational demands. In particular, a feature disentanglement module is introduced which is guided by two triplet losses: the first one, applied in a hyper-network fashion, encourages the removal of camera view information from the learned features; the second one enforces that similar driver actions have closer feature embeddings, while the representations of different actions are further apart. Moreover, this is the first work to conduct comprehensive experiments using the leave-one-camera-out strategy, accurately reflecting real-world scenarios where the recognition model is trained under certain conditions and tested under different ones. This rigorous setup ensures that our findings are both practical and applicable to real-world conditions.
@@ III-A Motivation
Image analysis-based DMSs must operate reliably in diverse real-world conditions, where camera placement and interior configuration vary significantly across vehicles [34, 8]. This cross-camera variability introduces a critical challenge: conventional supervised deep learning methods trained on a fixed viewpoint, often employing cross-entropy loss, struggle to learn robust representations capable of generalizing across different views and, therefore, their performance degrade when they are applied to unseen views. Most existing methods overlook this issue, focusing instead on maximizing accuracy within a fixed viewpoint – an unrealistic assumption for actual deployment. Moreover, DMSs must run efficiently on embedded hardware, where computational and memory resources are limited. Prior approaches typically either focus on model complexity reduction or on classification accuracy improvement, but rarely address both cross-camera robustness and resource efficiency.
In this work, we address both challenges jointly. We propose a learning strategy that explicitly disentangles driver action features from viewpoint-dependent cues, enabling the model to generalize across camera perspectives while maintaining a lightweight structure suitable for real-time applications. We introduce the feature disentanglement module, which is based on two complementary triplet losses to guide the network to learn robust, semantically meaningful representations. The feature disentanglement module uses the first triplet loss to train an action-independent view classifier and exploits it as a hyper-network to remove of camera view information from learned features. The second triplet loss is used to further reinforce viewpoint consistency across actions.
Our goal is not only to improve classification accuracy, but to ensure transferability and reliability in real deployment scenarios.
@@ III-B Overview
Given the factors mentioned above, we present the Driver Behavior Monitoring Network (DBMNet), illustrated in Figure 1. This method encodes an RGB image by leveraging the lightweight CNN architecture detailed in [8] as its backbone. The proposed feature disentanglement module aims to prioritize action encoding over camera-view information. This is achieved by linearly combining view queries with features estimated by the backbone. The resulting disentangled features are then forwarded to a linear layer for action classification. The model outputs two probability distributions: one for the driver’s action categories and another for the viewpoint. The viewpoint distribution acts as a hyper-network for the feature disentanglement module. The model is optimized through cross-entropy losses combined with triplet losses to assist in the learning of view-invariant action features.
Figure 1: The DBMNet pipeline. At inference time (a), the backbone encodes an RGB image into a feature vector $\mathbf{f}$. The feature disentanglement module then refines these features to discard view-related information while retaining action-related details using view queries, Q, together with the probability distribution over the views. The resulting features $\mathbf{\hat{f}}$ are mapped into a driver action by the action classifier. During training (b), alongside the cross-entropy loss for action and view prediction (respectively, $\mathcal{L}_{{\rm{ace}}}$ and $\mathcal{L}_{{\rm{vce}}}$), we employ the triplet loss ($\mathcal{L}_{{\rm{ac}}}$ and $\mathcal{L}_{{\rm{vc}}}$). This triplet loss helps to learn distinct action and view representations by processing three input images: an anchor image $a$, an image $sv$ depicting the same view but different action as $a$, and an image $sa$ with the same action but from a different view as $a$.
@@ III-C The proposed model
where $\mathbf{p}_{v}$ is the probability distribution over $V$ views estimated by the view classifier, which therefore acts as a hyper-network [45, 46]. The inner structure of this module might resemble memory banks, which are typically a learnable, or fixed, set for feature vectors or embeddings that the network can query and update: the idea of memory banks is to provide a module that can store feature representations so the model can retrieve and reuse them. Our feature disentanglement module instead learns a set of view queries that are independently multiplied for the feature vector coming from the backbone and linearly combined using as weights the probability distribution over the views that is estimated by the view classifier. As such, it also shares some similarities with the attention mechanism [47].
@@ III-D Training procedure
$\mathcal{L}=\mathcal{L}_{{\rm{ace}}}+\mathcal{L}_{{\rm{vce}}}+{\lambda_{\rm{ac}}}\mathcal{L}_{{\rm{ac}}}+{\lambda_{\rm{vc}}}\mathcal{L}_{{\rm{vc}}}.$ (5)
In the previous loss, there are two cross-entropy losses mainly dedicated to the optimization of linear layers for view ($\mathcal{L}_{{\rm{vce}}}$) and action ($\mathcal{L}_{{\rm{ace}}}$) categorization. Cross-entropy losses are only applied to the anchor image in this way:
$\mathcal{L}_{{\rm{vce}}}=-\sum_{m=1}^{V}y^{(m)}_{v}\mathrm{log}(p_{v}^{(m)}),$ (6)
$\mathcal{L}_{{\rm{ace}}}=-\sum_{m=1}^{A}y^{(m)}_{a}\mathrm{log}(p_{a}^{(m)}).$ (7)
@@
The previous two cross-entropy losses are complemented by two opposing triplet losses, whose contributions are controlled by the hyperparameters $\lambda_{\rm{ac}}$ and $\lambda_{\rm{vc}}$. Both losses rely on the triplet structure previously introduced. The difference lies in the role assumed by $sv$ and $sa$ in each loss.
In the action triplet loss (Eq. 8) the objective is to learn features for action encoding that are invariant regardless of the viewpoint. Hence, $sa$ serves as the positive sample, whereas $sv$ serves as the negative sample. This choice has the purpose of bringing closer the representations of $sa$ and $a$, which have the same action. Conversely, the model maximizes the distance between $a$ and $sv$, which represent different actions from the same viewpoint:
$\mathcal{L}_{{\rm{ac}}}=\sum_{n=1}^{N}[\delta+\mathcal{D}(\mathbf{\hat{f}}_{a},\mathbf{\hat{f}}_{sa})-\mathcal{D}(\mathbf{\hat{f}}_{a},\mathbf{\hat{f}}_{sv})],$ (8)
@@
For the view triplet loss, the positive and negative samples are reversed:
$\mathcal{L}_{{\rm{vc}}}=\sum_{n=1}^{N}[\delta+\mathcal{D}(\mathbf{f}_{a},\mathbf{f}_{sv})-\mathcal{D}(\mathbf{f}_{a},\mathbf{f}_{sa})].$ (9)
@@ 5 Texting (left) 11 Looking down 17 Adjusting radio
The driving sessions were recorded in both daytime and nighttime in five different vehicles, comprising two sedans, two SUVs, and one van. Within each vehicle, four cameras were strategically placed - on the left front, right front, and right side of the drivers. This arrangement served dual purposes. Firstly, having a dataset with multiple cameras aids in developing more robust models that can account for variations in camera perspectives, thus evaluating the models’ generalization across different cameras. Secondly, the multi-camera dataset offers an opportunity to enhance system performance by considering the content captured from multiple camera angles. Including nighttime recordings is also essential, as distraction-related behaviors – such as eye closure and yawning – are more frequent and harder to detect under low-light conditions. These recordings allow us to assess model robustness across varying lighting environments. Figure 2 reports samples from the four cameras in both day and night times.